OpenStudy (ryanchick):

In a school 50% of the students are younger than 10, 1/20 are 10 years old and 1/10 are older than 10 but younger than 12, the remaining 70 students are 12 years or older. How many students are 10 years old?

OpenStudy (anonymous):

Do you have any guesses or any options

OpenStudy (ryanchick):

nope

OpenStudy (anonymous):

I think it might it may be 4

OpenStudy (ryanchick):

i tried that its not right

OpenStudy (anonymous):

my second guess is 2

OpenStudy (ryanchick):

okay ill try that

OpenStudy (anonymous):

k

OpenStudy (ryanchick):

not right

OpenStudy (anonymous):

try 10

OpenStudy (ryanchick):

okay

OpenStudy (ryanchick):

not right

OpenStudy (ryanchick):

@jabez177 can you help

OpenStudy (aihberkhan):

Okay. 50% or \(0.50\) are younger than \(10\). \(0.10\) are older than \(10\) but they are also younger than \(12\) So, \(70\) students are \(12\) or older We need to find out how many are 10 years olf. How many are 10 yrs old? We would solve this equation: \(0.50+0.05+0.10=0.65=\) 65% So, 70 are 35% Now you would do, \(70/0.35=200\) in total So, \(200*0.05=10\) students ten years old

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) A medal would also be appreciated! Just click best response next to my name! Also, a fan would be appreciated as well! Just hover over my icon and click "Become A Fan". This will allow you to see every time I am online! :) If you see that I am online and need help with a question, just tag me in your question! @ryanchick

OpenStudy (anonymous):

@AihberKhan I tried that number already :P

OpenStudy (aihberkhan):

It should be 10... @DAVPAS1